OpenStudy (anonymous):

Jenny is eight years older than twice her cousin Sue’s age. The sum of their ages is less than 32. What is the greatest age that Sue could be? (Points : 1) A: 7 B: 8 C: 9 D: 10

OpenStudy (igreen):

@breathecarolina

OpenStudy (igreen):

If Sue was 7: Then Jenny would be 2 * 7 + 8 = 22 22 + 7 = 29.

OpenStudy (igreen):

Let's try 8: Then Jenny would be 2 * 8 + 8 = 38. So the oldest age Sue can be is 7.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @breathecarolina

OpenStudy (anonymous):

Thank you @iGreen

OpenStudy (igreen):

No problem.